# Spectrum Sharing in Satellite-Terrestrial Integrated Networks: Frameworks, Approaches, and Opportunities

### v2

@@ IV-B Interference Intensity between NTN and TN
Interference intensity refers to the power strength per unit of bandwidth caused by secondary users to primary users, which denotes the negative impact of secondary users when accessing the shared spectrum in STINs. SS brings additional interference into the system, as users receive interference from the TN nodes during NTN downlink transmission, and the TN nodes receive interference from the users during NTN uplink transmission. Hence, a smaller interference intensity implies a higher efficiency and capacity of the STINs. Since the total interference power grows as bandwidth increases, the STIN performance can be guaranteed by maintaining a satisfactory interference intensity.
@@ IV-D End-to-End Latency in STINs
End-to-end (E2E) latency refers to the time a user receives a data packet across the network once the packet is triggered. This metric is essential in the SS of STINs, as it reflects the timeliness and fairness of packet transmission. In STINs, secondary users may need to delay their transmissions if they cannot access the spectrum in the current time slot. Additionally, when dealing with satellite-to-ground communication links, it is essential to consider propagation delay, which can add several milliseconds to the E2E latency. Generally, a larger bandwidth indicates a higher data rate and eventually results in a smaller E2E latency.
@@ IV-E Energy Efficiency of NTN and/or TN
Energy efficiency (EE) refers to the number of bits a user can transmit by consuming per unit of power. As a key performance metric, EE is obtained by dividing data throughput by the amount of power consumption, which denotes energy utilization in transmitting signals. Since NTN nodes and users in STINs often face limited energy supply restrictions, it is crucial to promote their EE to transmit more data and provide better QoS. Meanwhile, increasing EE also enables us to better utilize scarce spectrum resources. Thus, a good guarantee of EE is essential for the overall system performance of STINs, and SS technique can naturally promote EE especially under large bandwidth conditions.
@@ V Case Study: SS in STINs Using Protection Zone-based Approach
FIGURE 5: System capacity and average data rate of STINs with SS and without SS. The total bandwidth is $300$ MHz, while the NTN bandwidth and the TN bandwidth are fixed to $120$ MHz and $180$ MHz, respectively, for STINs without SS. The BSs are randomly distributed in a $200$ km×$\times 200$ km remote area with a density of $\lambda_{\text{T}}=3\times 10^{-3}$/km2. Each BS has a service radius of $R_{\text{T}}=10$ km and transmit power of $P_{\text{T}}=46$ dBm. The pathloss exponent in TNs is $\alpha_{\text{T}}=3.5$. A typical TN user is assumed to be located at the center of the serving area, while typical NTN users are $50$ km away from the typical TN user. The noise power at the user side is $-110$ dBm. In NTNs, the satellites are distributed on an orbital sphere with a density of $\lambda_{\text{N}}=1\times 10^{-5}$/km2 at an altitude of $500$ km. The transmit power, main-lobe gain, side-lobe gain, and path loss exponent are $P_{\text{N}}=50$ dBm, $G_{\text{ml}}=30$ dBi, $G_{\text{sl}}=10$ dBi, and $\alpha_{\text{N}}\approx 2.0$, respectively.
@@
In this section, we quantitatively study the effectiveness of SS in STINs by adopting the protection zone-based approach. The simulation scenarios are first depicted in Fig. 4. The NTN downlink and TN downlink framework is considered, where the NTN users are the primary users and the TN users are the secondary users. A protection zone-based approach is adopted, so that the BSs are located at least $R_{\text{p}}$ away from NTN users. For the case of SS in Fig. 4(a), NTN users are primary users each having a protection zone with radius $R_{\text{p}}$. An NTN user will use the reserved spectrum if there is a BS inside the protection zone and will use the shared spectrum if there is no BS inside. If SS is not adopted in STINs as shown in Fig. 4(b), NTN users and TN users will experience interference from the NTN spectrum and TN spectrum, respectively. This case study evaluates the aforementioned networks using two key performance metrics. First, the average data rate is determined by multiplying the product of spectral efficiency and bandwidth. Second, the system capacity of the NTN (or TN) is assessed by multiplying the spectral efficiency by the satellite (or BS) density, the surface area of satellites (or BSs), and the available bandwidth.
@@
The simulation results are shown in Fig. 5, comparing the following three scenarios, i.e., Scenario 1: SS with the protection zone-based approach; Scenario 2: SS without the protection zone-based approach; Scenario 3: the setting without SS. Note that as reserved NTN bandwidth is considered in Scenario 2 and Scenario 3, the system capacity and average data rate of them remain almost the same. Specifically, in Fig. 5(a), the TN system capacity in Scenario 1 first surpasses that of Scenario 2 when $B_{\text{R}}$ is around $100$ MHz, but the latter subsequently exceeds it. After applying SS in Scenario 2, the excessive bandwidth can be utilized more efficiently, leading to an increase in sum and NTN system capacity compared to Scenario 3. The protection zone-based approach in Scenario 1 can utilize excessive reserved bandwidth and decrease co-frequency interference by setting an empirical protection radius $R_{\text{p}}$. If a BS is located within distance $R_{\text{p}}$, the NTN user will switch to the shared spectrum instead of the reserved spectrum. A gradual increase in both the sum and NTN system capacity can be observed.
@@
In Fig. 5(b), the average data rate of the users in STINs with SS is compared to that of the users in STINs without SS. Compared to Scenario 3, there will be higher data rate for NTN user but lower data rate to TN user in Scenario 2, which is mainly due to stronger co-frequency interference from satellite networks. In Scenario 1, with an increase of $B_{\text{R}}$, the data rate of the NTN user and the TN user using shared spectrum decreases because more bandwidth is allocated to the reserved spectrum. However, the data rate of the NTN user using reserved spectrum increases. Therefore, by selecting the appropriate reserved bandwidth according to implementation requirements, a win-win solution can be obtained by using SS with the protection zone design.
@@ VI Future Opportunities of SS in STINs
In this section, we briefly discuss future applications of SS in STINs. An overview of these potential applications is presented in Fig. 6, and detailed descriptions of each of them are given as follows.
@@ VI-B SS for Multi-Connectivity-Based Cooperative STINs
Multi-connectivity in STINs, which has been referred to in the 3GPP Technical Standard [14], indicates that a user can simultaneously connect to multiple network nodes, including BSs or access points in TNs and satellites in NTNs. With proper synchronization and mitigation of delay, phase misalignment, and Doppler shifts with algorithms [15], cooperative communication, e.g., relaying, joint transmission, or reception, can be utilized to connect different network tiers in STINs and promote more resilient and adaptable communications. However, the bandwidth allocated to a user, which may have multiple connection links, can be restricted due to the scarcity of available spectrum resources. Hence, the allocation of spectrum resources among multiple users, each possessing multiple links, presents significant research opportunities. In addition, the organization of cooperative network nodes, whether centralized or decentralized, requires further exploration and investigation.
@@ VI-C Multiple Access-Enabled Cognitive SS in STINs
Multiple access allows multiple users to be served simultaneously in the same frequency band. Using this characteristic, multiple access and cognitive SS can be connected by sensing the real-time radio environment. The cognitive radio can sense the surroundings and adjust the transmission parameters accordingly. With these sensing results, secondary users can assess the unutilized spectrum without interfering with primary users. In addition, this fusion also offers some other benefits, including increased network capacity, more efficient and flexible resource allocation, and decreased interference. By continuously monitoring the radio environment and adjusting transmission parameters, more reliable network performance can be ensured, and a more equitable distribution of spectrum resources among users can be promoted.
@@ VI-D SS for Different Satellite Constellations
As satellite constellations have expanded to mega size in recent years, the number of different satellite constellations built by different satellite companies is also increasing. Satellite constellations can also share their proprietary spectrum resources to improve the system’s capacity. Due to the mobile nature of the satellites, the NTN architecture is always dynamic. It changes rapidly, leading to the significant challenge of maintaining reliable and high-quality services under the restriction of limited and congested spectrum resources. Different satellite constellations face their transmission conditions as service providers. They require compatible spectrum access strategies. SS techniques enable the high-efficiency utilization of spectrum resources while introducing controllable interference. Meanwhile, to meet the constellation dynamics, the designed SS strategy should also be adaptive to available satellites and spectrum resources from time to time in providing wireless services.
@@ VII Conclusions
This article investigated the frameworks, approaches, performance metrics, and future research opportunities for SS in STINs. We introduced four general SS frameworks in STINs, including different uplink and downlink combinations for NTN and TN, and discussed their characteristics. Then, state-of-the-art approaches for SS to reuse spectrum resources in space, time, and frequency domains in STINs were comprehensively reviewed. More importantly, related metrics were presented to study the SS techniques further and evaluate the performance of the system and the individual user. A case study on the protection zone-based approach for SS in STINs was carried out, and the simulation results showed that a win-win solution could be obtained for improved system capacity. Finally, we provided promising opportunities and insightful future research on SS in STINs.
@@
Bodong Shang (bdshang@eitech.edu.cn) received his M.S. degree from Xidian University, China, and his Ph.D. degree from the Department of Electrical and Computer Engineering at Virginia Tech, Blacksburg, USA, in 2021, and he was a Postdoctoral Research Associate at Carnegie Mellon University, Pittsburgh, USA. He held a research internship position at Nokia Bell Labs, USA. He is an Assistant Professor at Eastern Institute for Advanced Study, Eastern Institute of Technology (EIT), Ningbo, China. His research interests include space-air-ground-sea integrated networks, non-terrestrial networks, and space information networks.
@@
Chungang Yang (guideyang2050@163.com) is a full professor at Xidian University, leading the GUIDE: Game, Utility, Artificial Intelligent Design for Emerging Communications research team. His research interests include artificial intelligence 6G wireless mobile networks, intent-driven networks, space-terrestrial networks, and game theory for emerging communication networks.
Chao Ren (chaoren@ustb.edu) is an Associate Professor and Deputy Director with the University of Science and Technology in Beijing, China. He is an Editorial Board Member for the Journal of Communications and Frontiers in Computer Science, recognized with an Outstanding Editor Award. He has been a Guest Editor for Space-Integrated-Ground Information Networks and Information Security Research, Co-Chair of the UCOM 2023-2025 Workshop, and Technical Program Committee member for over 30 international conferences in communications.
Haijun Zhang (haijunzhang@ieee.org) is a full professor and dean at the University of Science and Technology in Beijing, China. He serves/served as an Editor of IEEE Transactions on Information Forensics and Security, IEEE Transactions on Communications, IEEE Transactions on Network Science and Engineering, and IEEE Transactions on Vehicular Technology. He received the IEEE CSIM Technical Committee Best Journal Paper Award in 2018, the IEEE ComSoc Young Author Best Paper Award in 2017, and the IEEE ComSoc Asia-Pacific Best Young Researcher Award in 2019. He is a Fellow of IEEE.

### v3

@@ IV-B Interference Intensity between NTN and TN
Interference intensity refers to the power strength per unit of bandwidth caused by secondary users to primary users, which denotes the negative impact of secondary users when accessing the shared spectrum in STINs. SS introduces additional interference into the system, as users experience interference from TN nodes during NTN downlink transmission, and TN nodes receive interference from users during NTN uplink transmission. Hence, a smaller interference intensity implies a higher efficiency and capacity of the STINs. Since the total interference power increases as the bandwidth grows, the STIN performance can be guaranteed by maintaining a satisfactory interference intensity.
@@ IV-D End-to-End Latency in STINs
End-to-end (E2E) latency refers to the time it takes for a user to receive a data packet across the network, from the moment the packet is triggered. This metric is essential in the SS of STINs, as it reflects the timeliness and fairness of packet transmission. In STINs, secondary users may need to delay their transmissions if they cannot access the spectrum in the current time slot. Additionally, when dealing with satellite-to-ground communication links, it is essential to consider propagation delay, which can add several milliseconds to the E2E latency. Generally, a larger bandwidth indicates a higher data rate, which ultimately results in a smaller E2E latency.
@@ IV-E Energy Efficiency of NTN and/or TN
Energy efficiency (EE) refers to the number of bits a user can transmit by consuming per unit of power. As a key performance metric, EE is calculated by dividing the data throughput by the power consumption, which represents the energy utilization in transmitting signals. Since NTN nodes and users in STINs often face limited energy supply restrictions, it is crucial to promote their EE to transmit more data and provide better QoS. Meanwhile, increasing EE also enables us to better utilize scarce spectrum resources. Thus, a good guarantee of EE is essential for the overall system performance of STINs, and the SS technique can naturally promote EE, especially under large bandwidth conditions.
@@ V Case Study: SS in STINs Using Protection Zone-based Approach
FIGURE 5: System capacity and average data rate of STINs with SS and without SS. The total bandwidth is $300$ MHz, while the NTN bandwidth and the TN bandwidth are fixed to $120$ MHz and $180$ MHz, respectively, for STINs without SS. The BSs are randomly distributed in a $200$ $\times 200 km$ remote area with a density of $\lambda_{\text{T}}=3\times 10^{-3}$/km2. Each BS has a service radius of $R_{\text{T}}=10$ km and transmit power of $P_{\text{T}}=46$ dBm. The pathloss exponent in TNs is $\alpha_{\text{T}}=3.5$. A typical TN user is assumed to be located at the center of the serving area, while typical NTN users are $50$ km away from the typical TN user. The noise power at the user side is 110-110 dBm. In NTNs, the satellites are distributed on an orbital sphere with a density of $\lambda_{\text{N}}=1\times 10^{-5}$/km2 at an altitude of $500$ km. The transmit power, main-lobe gain, side-lobe gain, and path loss exponent are $P_{\text{N}}=50$ dBm, $G_{\text{ml}}=30$ dBi, $G_{\text{sl}}=10$ dBi, and $\alpha_{\text{N}}\approx 2.0$, respectively.
@@
In this section, we quantitatively study the effectiveness of SS in STINs by adopting the protection zone-based approach. The simulation scenarios are first depicted in Fig. 4. The NTN downlink and TN downlink framework is considered, where the NTN users are the primary users and the TN users are the secondary users. A protection zone-based approach is adopted, so that the BSs are located at least $R_{\text{p}}$ away from NTN users. For the case of SS in Fig. 4(a), NTN users are primary users, each having a protection zone with radius $R_{\text{p}}$. An NTN user will use the reserved spectrum if there is a BS inside the protection zone and will use the shared spectrum if there is no BS inside. If SS is not adopted in STINs as shown in Fig. 4(b), NTN users and TN users will experience interference from the NTN spectrum and TN spectrum, respectively. This case study evaluates the aforementioned networks using two key performance metrics. First, the average data rate is determined by multiplying the product of spectral efficiency and bandwidth. Second, the system capacity of the NTN (or TN) is assessed by multiplying the spectral efficiency by the satellite (or BS) density, the surface area of satellites (or BSs), and the available bandwidth.
@@
The simulation results are shown in Fig. 5, comparing the following three scenarios, i.e., Scenario 1: SS with the protection zone-based approach; Scenario 2: SS without the protection zone-based approach; Scenario 3: the setting without SS. Note that, as reserved NTN bandwidth is considered in Scenarios 2 and 3, the system capacity and average data rate of these scenarios remain almost the same. Specifically, in Fig. 5(a), the TN system capacity in Scenario 1 first surpasses that of Scenario 2 when $B_{\text{R}}$ is around $100$ MHz, but the latter subsequently exceeds it. After applying SS in Scenario 2, the excessive bandwidth can be utilized more efficiently, leading to an increase in the sum and NTN system capacity compared to Scenario 3. The protection zone-based approach in Scenario 1 can utilize excessive reserved bandwidth and decrease co-frequency interference by setting an empirical protection radius $R_{\text{p}}$. If a BS is located within distance $R_{\text{p}}$, the NTN user will switch to the shared spectrum instead of the reserved spectrum. A gradual increase in both the sum and the NTN system capacity can be observed.
@@
In Fig. 5(b), the average data rate of the users in STINs with SS is compared to that of the users in STINs without SS. Compared to Scenario 3, the NTN user will have a higher data rate, but the TN user will have a lower data rate in Scenario 2, primarily due to more substantial co-frequency interference from satellite networks. In Scenario 1, with an increase of $B_{\text{R}}$, the data rate of the NTN user and the TN user using shared spectrum decreases because more bandwidth is allocated to the reserved spectrum. However, the data rate of the NTN user using the reserved spectrum increases. Therefore, by selecting the appropriate reserved bandwidth according to implementation requirements, a win-win solution can be obtained by using SS with the protection zone design.
@@ VI Future Opportunities of SS in STINs
In this section, we briefly discuss future applications of SS in STINs. An overview of these potential applications is presented in Fig. 6, and detailed descriptions of each are provided below.
@@ VI-B SS for Multi-Connectivity-Based Cooperative STINs
Multi-connectivity in STINs, as referred to in the 3GPP Technical Standard [14], enables a user to simultaneously connect to multiple network nodes, including BSs or access points in TNs and satellites in NTNs. With proper synchronization and mitigation of delay, phase misalignment, and Doppler shifts using algorithms [15], cooperative communication, such as relaying, joint transmission, or reception, can be utilized to connect different network tiers in STINs and promote more resilient and adaptable communications.
However, the bandwidth allocated to a user, which may have multiple connection links, can be restricted due to the scarcity of available spectrum resources. Hence, the allocation of spectrum resources among multiple users, each possessing multiple links, presents significant research opportunities. Additionally, the organization of cooperative network nodes, whether centralized or decentralized, warrants further exploration and investigation.
@@ VI-C Multiple Access-Enabled Cognitive SS in STINs
Multiple access allows multiple users to share the same spectrum simultaneously. The cognitive radio can sense the real-time radio environment and adjust the transmission parameters accordingly. With these sensing results, secondary users can assess the unutilized spectrum without interfering with primary users. For instance, non-orthogonal multiple access (NOMA) allows users to share the same frequency resources with different power levels. This can be paired with spectrum sensing to adjust user clustering and power allocation based on the levels of interference. Rate-splitting multiple access (RSMA) can flexibly split messages into common and private parts, which can be adapted to the variable interference landscape in shared spectrum environments.
Using these multiple access schemes, STINs can not only increase spectrum efficiency but also support user fairness and low-latency requirements under dynamic spectrum availability. Future research can focus on jointly optimizing spectrum sensing, access control, and user scheduling for each multiple-access scheme.
@@ VI-D SS for Different Satellite Constellations
As satellite constellations have expanded to mega size in recent years, the number of different satellite constellations built by different satellite companies is also increasing. Satellite constellations can also share their proprietary spectrum resources to improve the system’s capacity. Due to the mobile nature of the satellites, the NTN architecture is always dynamic. It changes rapidly, posing a significant challenge to maintaining reliable and high-quality services under the constraints of limited and congested spectrum resources. Different satellite constellations face different transmission conditions as service providers. They require compatible spectrum access strategies. SS techniques enable the high-efficiency utilization of spectrum resources while introducing controllable interference. Meanwhile, to meet the constellation dynamics, the designed SS strategy should also be adaptive to the availability of satellites and spectrum resources over time in providing wireless services.
@@ VII Conclusions
This article investigated the frameworks, approaches, performance metrics, and future research opportunities for SS in STINs. We introduced four general SS frameworks in STINs, including different uplink and downlink combinations for NTN and TN, and discussed their characteristics. Then, state-of-the-art approaches for SS to reuse spectrum resources in space, time, and frequency domains in STINs were comprehensively reviewed. More importantly, related metrics were presented to study the SS techniques further and evaluate the system’s and individual users’ performance. A case study on the protection zone-based approach for SS in STINs was conducted, and the simulation results demonstrated that a win-win solution could be achieved for enhanced system capacity. Finally, we provided promising opportunities and insightful future research on SS in STINs.
@@
Bodong Shang (bdshang@eitech.edu.cn) received his Ph.D. degree from the Department of Electrical and Computer Engineering at Virginia Tech, Blacksburg, USA. He was a Postdoctoral Research Associate at Carnegie Mellon University, Pittsburgh, USA. He held a research internship position at Nokia Bell Labs, USA. He is an Assistant Professor at Eastern Institute for Advanced Study, Eastern Institute of Technology (EIT), Ningbo, China. His research interests include space-air-ground-sea integrated networks, non-terrestrial networks, and space information networks.
@@
Chungang Yang (guideyang2050@163.com) is a Professor at Xidian University, leading the GUIDE: Game, Utility, Artificial Intelligent Design for Emerging Communications research team. His research interests include artificial intelligence, 6G wireless mobile networks, intent-driven networks, space-terrestrial networks, and game theory for emerging communication networks.
Chao Ren (chaoren@ustb.edu.cn) is an Associate Professor and Deputy Director with the University of Science and Technology in Beijing, China. He is an Editorial Board Member for the Journal of Communications and Frontiers in Computer Science, recognized with an Outstanding Editor Award. He has been a Guest Editor for Space-Integrated-Ground Information Networks and Information Security Research, Co-Chair of the UCOM 2023-2025 Workshop, and Technical Program Committee member for over 30 international conferences in communications.
Haijun Zhang (haijunzhang@ieee.org) is a Professor and dean at the University of Science and Technology in Beijing, China. He serves/served as an Editor of IEEE Transactions on Information Forensics and Security, IEEE Transactions on Communications, IEEE Transactions on Network Science and Engineering, and IEEE Transactions on Vehicular Technology. He received the IEEE CSIM Technical Committee Best Journal Paper Award in 2018, the IEEE ComSoc Young Author Best Paper Award in 2017, and the IEEE ComSoc Asia-Pacific Best Young Researcher Award in 2019. He is a Fellow of IEEE.